Q If 7 th term of a H.P is 8 and 8 th term is 7, then its 15 th term is?

Q If 7th term of a H.P is 8 and 8th term is 7, then its 15th term is?

Grade:11

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given 7thterm of a HP is (1/a+6d)=8
8thtermis (1/a+7d)=7
from above equations we get d=(1/7)-(1/8)=1/56
and a=(1/8)-(6/56)=1/56
15thterm of HP=1/(a+14d)=56/15
Thanks and Regards
Shaik Aasif
askIITians faculty